# Scalable registration of single quantum emitters within solid immersion lenses through femtosecond laser writing

### v1

@@ IV Conclusions
In this paper, we have demonstrated marker-free registration of single quantum emitters at the center of SILs, by femtosecond laser writing. The microlens both lower the laser dose required to generate the emitter and aid in positioning it near the center of the structure, with an accuracy of within in the y-direction and in the x-direction.
There remains an open question regarding the nature and diversity of created emitters and their performance for quantum technology applications. The formation of carbon-related defects, such as carbon vacancies, in SiC is known to be energetically quite favourable, suggesting that the laser irradiation might be creating other carbon-related defects in the immediate vicinity of the optically active VSi. ODMR properties have been measured, and none provided identifiable peaks. A further study after an anneal of the sample demonstrated that there are options to improve the spin properties of the defects. However, the yield of these results was low. Future studies will identify these defects and characterize their properties as a function of the laser treatment parameters.
In conclusion, this paper demonstrates the generation of single quantum emitters registered directly into microlens photonic structures in SiC, which can open new avenues for the scalable integration of spin-photon interfaces into quantum devices.
@@ acknowledgments
We express our gratitude to Nguyen Tien Son and Ivan G. Ivanov for valuable discussions. We also thank Pasquale Cilibrizzi for his helpful guidance during the low-temperature measurements. This work is funded by the Royal Academy of Engineering (RF2122-21-129), the Engineering and Physical Sciences Research Council (EP/S000550/1, EP/V053779/1, EP/Z533208/1, EP/Z533191/1, EP/W025256/1), the European Commission (QuanTELCO, grant agreement No 862721), the Leverhulme Trust (RPG-2019-388) and Euramet (23NRM04 NoQTeS).
@@
S0.1 Laser writing system
The laser writing system employed in this work is consistent with that described in a previous study [51]. The system comprises a Spectra Physics Mai Tai laser and a Spectra Physics Solstice amplifier, operating at a wavelength of 790 nm with a maximum repetition rate of 1 kHz. Pulse energy is regulated via a waveplate in combination with a Glan-laser polarizer. The beam is subsequently expanded before reaching the spatial light modulator (SLM, Hamamatsu Photonics X10468-02). A dichroic mirror directs the fabrication laser toward the sample while transmitting both the excitation and PL signals from an integrated room-temperature confocal system, featuring a 532 nm CW laser as the excitation source. Schematics of the laser writing system and confocal microscope are provided in Fig. S2 and Fig. S2, respectively.
S0.2 Description of the circle-fitting procedure
Here we describe the procedure to assess the position of the created emitter with respect to the SIL, based on PL maps with a stepsize of 0.13 m. First of all, we determine the SIL circular edge to retrieve its center. We used a data analysis application (ImageJ) to determine the SIL circular edges, using three different methods to verify the reproducibility of the outcomes. First, we use the fit circle tool to fit an exact circle around the SIL; second, we use the ellipse tool to fit precisely the SIL edge. Lastly, we used the profile tool to measure the changes in PL, which are expected to be high at the SIL edge as the surface of SIL itself fluoresces. We then fit a Gaussian to each of these PL peaks to pinpoint a precise position for the edges. Figure S3 shows a comparison between the three methods for the same SIL. The ellipse method in orange, the circle method in yellow and the blue lines indicate where the profiles would be taken to find the SIL edge as the PL counts rise relative to the background within the SIL.
Similarly, we compare results from two methods to determine the emitter position. First, we use the fit circle tool in ImageJ, we fit a circle around the emitter confocal spot, and retrieve its center. Second, we fit a Gaussian to the emitter confocal spot. Figure S3 shows the typical positions for these methods, where the blue lines again show the profile positions and the green shows the circle that would fit around the defect to determine the central point to be taken as the defect position.
@@
S0.4 Optical Spectra of Emitters
@@
The analysis uncovered a broad distribution of zero-phonon lines (ZPL) within the 858nm-985nm spectral region. Out of 39 examined defect centers, six exhibited spectral lines typical of silicon-vacancy (VSi) centers: one V1’ center (ZPL at 858nm), four V1 centres (h-site, ZPL at 861nm), and one V2 center (k-site, ZPL at 916nm). Other defect centers showed ZPLs at various wavelengths within the 858-985nm range, which do not correspond to previously reported unidentified lines in this range but align with the theoretically predicted range of VSi centers modified by nearby carbon anti-sites.t
@@
S0.5 ODMR after Anneal
An anneal at 600 degrees was conducted for 30 minutes on the sample in vacuum at mbar to investigate the stability of the observed defects, and whether they could be converted to silicon vacancies. We re-characterised 5 SILs, which originally showed no ODMR. One evidenced an ODMR peak after annealing, with 0.48% contrast and 22.36 MHz linewidth (figure S6).
The ODMR measurement is performed by focusing an off-resonant cw-laser (730 nm) on the a-plane side of the 4H-SiC sample inside the cryostat (Attocube Attodry 800) at 4K with a high-NA objective (Zeiss Epiplan-Neofluar 100x, NA 0.9). The spin state manipulation of the color center is orchestrated with the help of an Arbitrary Waveform Generator (QM-OPX) and a copper wire (50 diameter) running on top of the sample. The AWG feeds a radio-frequency signal with a power of -27dBm, amplified by a 44 dB amplifier (LZY-22+ Mini-circuits) and fed to the microwave antenna wire running over the sample. The manipulation of the spin states is realized once the frequency of microwave matches with the zero field splitting of ground state spin ( 70 MHz for a color center at = 0G). The difference in spin state population while sweeping the MW frequency is seen through the photoluminescence emitted by the V2 VSi center which is filtered using a dichroic mirror (Semrock FF925-Di01) and along pass filter (FELH950 Thorlabs). The emission is readout through a super conducting nanowire single photon detector (Single Quantum).

### v2

@@ Abstract
In this work, we have demonstrated marker-free registration of single quantum emitters at the center of SILs by femtosecond laser writing. The SILs lower the laser dose required to generate emitters and aid in positioning them near the center of the structure, with an accuracy of in the y-direction and in the x-direction. Our approach, based on direct alignment of the laser-writing beam focus to the center of the SIL, achieves sub-diffraction-limit placement accuracy within the SIL to maximize optical collection enhancement. In future applications, better positioning (potentially nm) could be achieved through use of alignment markers.
There remains an open question regarding the nature and diversity of created emitters and their performance for quantum technology applications. In our postannealing study, controlled thermal treatment revealed opportunities to enhance the spin coherence of the generated defects; however, the yield of highcoherence centers remained low. Future investigations could systematically identify generated defect species and map their optical and spincoherence properties as functions of laser irradiation parameters and thermal annealing protocols. Furthermore, deterministic generation of V centers can be further pursued by integrating laserannealing sequences with in-situ photoluminescence monitoring.
Supporting Information: Experimental details of ODMR measurements, optical spectra, and XZ profiles, descriptions of the experimental setups, fitting procedure details, and sample preparation methods.
@@ acknowledgments
We express our gratitude to Nguyen Tien Son and Ivan G. Ivanov for valuable discussions. This work is funded by the Royal Academy of Engineering (RF2122-21-129), the Engineering and Physical Sciences Research Council (EP/S000550/1, EP/V053779/1, EP/Z533208/1, EP/Z533191/1, EP/W025256/1), the European Commission (QuanTELCO, grant agreement No 862721), the Leverhulme Trust (RPG-2019-388). It is further supported by the project 23NRM04 NoQTeS, which has received funding from the European Partnership on Metrology, co-financed from the European Union’s Horizon Europe Research and Innovation Programme and by the Participating State.
@@
S1 Sample Preparation
The experiment was conducted using commercial 4H-SiC material (Xiamen PowerWay©) diced into chips.With substrate and epilayer thickness and respectively, and residual n-doping level of . The material was diced into chips. Arrays of hemispherical SILs with nominal radius were fabricated on these chips using the grayscale hard-mask lithography process set out in our previous work [26].
S2 Laser writing system
The laser writing system employed in this work is consistent with that described in a previous study [54]. The system comprises a Spectra Physics Mai Tai laser and a Spectra Physics Solstice amplifier, operating at a wavelength of 790 nm with a maximum repetition rate of 1 kHz. Pulse energy is regulated via a waveplate in combination with a Glan-laser polarizer. The beam is subsequently expanded before reaching the spatial light modulator (SLM, Hamamatsu Photonics X10468-02). A dichroic mirror directs the fabrication laser toward the sample while transmitting both the excitation and PL signals from an integrated room-temperature confocal system, featuring a 532 nm CW laser as the excitation source. Laser fabrication and PL characterization employed an Olympus PlanApo 60× oilimmersion objective (NA = 1.4) with immersion oil of refractive index of n = 1.5. Emission under 532 nm excitation was spectrally filtered to the 600–800 nm window using a bandpass filter and a notch filter. Schematics of the laser writing system and confocal microscope are provided in Fig. S1 and Fig. S2, respectively.
S3 Description of the circle-fitting procedure
Here we describe the procedure to assess the position of the created emitter with respect to the SIL, based on PL maps with a stepsize of 0.13 m. First of all, we determine the SIL circular edge to retrieve its center. We used a data analysis application (ImageJ) to determine the SIL circular edges, using three different methods to verify the reproducibility of the outcomes. First, we use the fit circle tool to fit an exact circle around the SIL; second, we use the ellipse tool to fit precisely the SIL edge. Lastly, we used the profile tool to measure the changes in PL, which are expected to be high at the SIL edge as the surface of SIL itself fluoresces. We then fit a Gaussian to each of these PL peaks to pinpoint a precise position for the edges. Fig. S3 shows a comparison between the three methods for the same SIL. The ellipse method in orange, the circle method in yellow and the blue lines indicate where the profiles would be taken to find the SIL edge as the PL counts rise relative to the background within the SIL.
Similarly, we compare results from two methods to determine the emitter position. First, we use the fit circle tool in ImageJ, we fit a circle around the emitter confocal spot, and retrieve its center. Second, we fit a Gaussian to the emitter confocal spot. Fig. S3 shows the typical positions for these methods, where the blue lines again show the profile positions and the green shows the circle that would fit around the defect to determine the central point to be taken as the defect position.
@@
S5 Optical Spectra of Emitters
@@
The analysis uncovered a broad distribution of zero-phonon lines (ZPL) within the 858-985 nm spectral region. Out of 39 examined defect centers, six exhibited spectral lines typical of silicon-vacancy (V) centers: one V1’ center (ZPL at 858 nm), four V1 centres (h-site, ZPL at 861 nm), and one V2 center (k-site, ZPL at 916 nm). Other defect centers showed ZPLs at various wavelengths within the 858-985 nm range, which do not correspond to previously reported unidentified lines in this range but align with the theoretically predicted range of V centers modified by nearby carbon anti-sites.
@@
S6 ODMR after Anneal
An anneal at 600 degrees was conducted for 30 minutes on the sample in vacuum at mbar to investigate the stability of the observed defects, and whether they could be converted to silicon vacancies. We re-characterised 5 SILs, which originally showed no ODMR. One evidenced an ODMR peak after annealing, with 0.48% contrast and 22.36 MHz linewidth (Fig. S6).
The ODMR measurement is performed by focusing an off-resonant cw-laser (730 nm) on the a-plane side of the 4H-SiC sample inside the cryostat (Attocube Attodry 800) at 4K with a high-NA objective (Zeiss Epiplan-Neofluar 100x, NA 0.9). The spin state manipulation of the color center is orchestrated with the help of an Arbitrary Waveform Generator (QM-OPX) and a copper wire (50 diameter) running on top of the sample. The AWG feeds a radio-frequency signal with a power of -27dBm, amplified by a 44 dB amplifier (LZY-22+ Mini-circuits) and fed to the microwave antenna wire running over the sample. The manipulation of the spin states is realized once the frequency of microwave matches with the zero field splitting of ground state spin ( 70 MHz for a color center at = 0G). The difference in spin state population while sweeping the MW frequency is seen through the photoluminescence emitted by the V2 V center which is filtered using a dichroic mirror (Semrock FF925-Di01) and along pass filter (FELH950 Thorlabs). The emission is readout through a super conducting nanowire single photon detector (Single Quantum).
The filtering during spectral acquisition was carried out using a SEMROCK long-pass filter in the transmission direction, oriented at approximately to achieve an effective cutoff near 900 nm. This configuration was chosen to remove stray light from the 730 nm excitation laser, which has a mild band extending to 1000 nm. The 900 nm filter allows transmission of zero-phonon lines (ZPLs), including the center at 916 nm. To enhance clarity for the reader, cosmic ray artefacts were removed during post-processing.